# Constrained mean-field control with singular control: Existence, stochastic maximum principle and constrained FBSDE

### v2

@@ 4 SMP by the Lagrange Multipliers Method
Let us first introduce some basic spaces of functions and processes. Let $\mathcal{X}$ be the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. Let $\mathbb{L}_{\mathbb{F}}^{2}$ be the set of $\mathbb{F}$-adapted processes $\beta=(\beta_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{l}$ such that $\mathbb{E}[\int_{0}^{T}|\beta_{t}|^{2}\mathrm{d}t]<\infty$. As a result, $\mathbb{L}_{\mathbb{F}}^{2}$ is a Hilbert space equipped with the inner product $\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}=\mathbb{E}[\int_{0}^{T}\langle\beta^{1}_{t},\beta^{2}_{t}\rangle\mathrm{d}t]$ for $\beta^{1},\beta^{2}\in\mathbb{L}_{\mathbb{F}}^{2}$. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2}})$. Define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
@@
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $\eta^{1}=(\eta_{t}^{1})_{t\in[0,T]}$ and $\eta^{2}=(\eta_{t}^{2})_{t\in[0,T]}$ take values respectively in $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$, which are $\mathbb{F}$-adapted processes such that $\mathbb{E}[\int_{0}^{T}(|\eta_{t}^{1}|^{2}+|\eta_{t}^{2}|^{2})\mathrm{d}t]<\infty$. Thus, $\mathbb{M}^{2}$ is a Hilbert space under the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by
@@
To adopt the methodology introduced in [6], we propose to embed the state process into a suitable Banach space. Without loss of generality, we can assume that for any singular control $\zeta\in\mathcal{A}^{k}[0,T]$, the associated radon measures $\mu_{i}$ for $i\in I$ of the trajectory $t\mapsto\zeta_{t}(\omega)$ is singular w.r.t. Lebesgue measure. Otherwise, we can decompose $\mu_{i}$ using the Lebesgue decomposition w.r.t. the Lebesgue measure and absorb the absolutely continuous (regular) term to the drift $b$. Consider the following space:
$\mathscr{X}:=\Big\{X=(X_{t})_{t\in[0,T]};~X\text{ is }\mathbb{F}\text{-adapted process with $X_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ and }$ (14)
$~X_{\cdot}=X_{0}+\int_{0}^{\cdot}X_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}X_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}X_{s}^{3}~\text{with}~(X^{1},X^{2},X^{3})\in\mathbb{L}_{\mathbb{F}}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}\times\mathcal{A}^{k}[0,T]\Big\}.$
@@
$\|X\|_{\mathscr{X}}:=\|X^{1}\|_{\mathbb{L}_{\mathbb{F}}^{2}}+\|X^{2}\|_{\mathbb{L}_{\mathbb{F}}^{2}}+\sqrt{\mathbb{E}\left[\left|\int_{0}^{T}G(s)\mathrm{d}X_{s}^{3}\right|^{2}\right]},~~\forall X\in\mathscr{X}.$
By the assumption that, for any $X\in\mathscr{X}$, its 3rd component $X^{3}$ of the decomposition in $\mathscr{X}$ is singular w.r.t. to Lebesgue measure, one can conclude that $\mathscr{X}$ is isometric to the product space $\mathbb{M}^{2}\times\mathbb{G}^{2}$ under the product norm, where $(\mathbb{G}^{2},\|\cdot\|_{\mathbb{G}^{2}})$ is a Banach space defined by
$\mathbb{G}^{2}:=\left\{S=(S_{t})_{t\in[0,T]};~S_{\cdot}=\int_{0}^{\cdot}G(s)\mathrm{d}\eta_{s}~\text{for}~\eta=(\eta_{t})_{t\in[0,T]}\in\mathcal{A}^{k}[0,T]\right\},$
$\|S\|_{\mathbb{G}^{2}}:=\sqrt{\mathbb{E}\left[\left|\int_{0}^{T}G(s)\mathrm{d}\eta_{s}\right|^{2}\right]}.$
Since $\int_{0}^{\cdot}G(s)\mathrm{d}\eta_{s}\in(\mathcal{M}[0,T])^{n}$ for $\eta\in\mathcal{A}^{k}[0,T]$, we can identify $\mathbb{G}^{2}$ as a subset of $L^{2}(\Omega;(\mathcal{M}[0,T])^{n})$ with $(\mathcal{M}[0,T])^{n}$ equipped with the norm $\|\mu\|_{(\mathcal{M}[0,T])^{n}}=\sqrt{\sum_{i=1}^{n}\|\mu_{i}\|^{2}_{\mathrm{TV}}}$. Thus, $\mathscr{X}$ can be viewed as a subset of $\mathbb{T}^{2}:=\mathbb{M}^{2}\times L^{2}(\Omega;(\mathbb{M}[0,T])^{n})$ under the product norm. For any $\lambda=(\lambda^{1},\lambda^{2})\in\mathbb{T}^{2,*}\simeq\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{n})$, the dual pair of $X=L+S\in\mathscr{X}$ for $(L,S)\in\mathbb{M}^{2}\times\mathbb{G}^{2}$ is defined by
$\langle X,\lambda\rangle_{\mathbb{T}^{2},\mathbb{T}^{2,*}}=\langle L,\lambda^{1}\rangle_{\mathbb{M}^{2}}+\langle S,\lambda^{2}\rangle_{L^{2}(\Omega;(\mathbb{M}[0,T])^{k}),L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{k})}.$
@@
$\mathtt{X}:=\mathbb{T}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}\times L^{2}(\Omega;(\mathbb{M}[0,T])^{n}),\quad\mathtt{Y}_{i}:=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},~\forall i\in I,\quad\mathtt{H}:=\mathbb{T}^{2}.$ (15)
@@
$J(X,\alpha,\zeta)$ $:=\mathbb{E}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}(X_{t}),\alpha_{t})\mathrm{d}t+\int_{0}^{T}c(t)\mathrm{d}\zeta_{t}\right],$
@@
$\begin{cases}\displaystyle\inf~J(X,\alpha,\zeta)\\[3.99994pt] \displaystyle\mathrm{s.t.}~~(X,\alpha,\zeta)\in C:=\mathscr{X}\times\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T],\\[3.99994pt] \displaystyle~~~~~~~\Phi_{i}(X,\alpha,\zeta)\leq_{K}0,~~\forall i\in I,\\[3.99994pt] \displaystyle\displaystyle~~~~~~~F(X,\alpha,\zeta)=0,\end{cases}$ (17)
@@
By using Lemma 3.11 in [6], the FJ optimality condition holds for the constrained optimization problem (17) in the following sense.
@@
Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal pair to problem (17). Then, there exist multipliers $r_{0}\geq 0$, $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ and $\lambda=(\lambda_{1},\lambda_{2})\in\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T],\mathbb{R})^{**})^{n})$, which are not all zeros, such that
@@
$\qquad\quad-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)\right\rangle_{\mathtt{H}^{*},\mathtt{H}}~~\text{\rm in}~\mathbb{L}_{\mathbb{F}}^{2},$
@@
where $\xi=(\xi_{1},\xi_{2},\xi_{3})\in N_{C}(\widehat{X},\widehat{\alpha},\widehat{\zeta})\subset\mathtt{X}^{*}$ and $\|\cdot\|_{\rm TV}$ is the total variation norm on the space of (signed) Radon measures.
@@
Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal solution to problem (3). There exist some $r_{0}\geq 0$ and $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ such that
@@
Furthermore, the stochastic (first-order) minimum condition holds that, $\mathrm{d}t\otimes\mathrm{d}\mathbb{P}$-a.s.
@@
$Y_{t}$ $=\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
$\langle\xi_{1},\widehat{X}-X\rangle_{\mathbb{T}^{2}}+\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}+\langle\xi_{3},\zeta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}\geq 0.$
Taking $\alpha=\widehat{\alpha}$ and $\zeta=\widehat{\zeta}$, we have $\xi_{1}=0$, $\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}\geq 0$ and $\langle\xi_{3},\zeta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathbb{M}[0,T])^{n}}$ $\geq 0$ similar as [6] (there is a slight change in the proof of Lemma 3.10 to fit into our setting here). Then, we can conclude from the structure of Lions derivative (c.f. Section 5.2.1 in [10]) and Lemma 4.1 that
$0=r_{0}\mathbb{E}\left[\int_{0}^{T}\langle\nabla_{x}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t}),L_{t}\rangle\mathrm{d}t+\mathbb{E}^{\prime}\left[\int_{0}^{T}\langle\partial_{\mu}f(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime}),L_{t}^{\prime}\rangle\mathrm{d}t\right]\right]$
$-\sum_{i\in I}\mathbb{E}\left[\int_{0}^{T}\left\langle\nabla_{x}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})+\mathbb{E}^{\prime}\left[\partial_{\mu}\phi^{i}(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})\right],L_{t}^{\prime}\right\rangle\eta^{i}_{t}\mathrm{d}t\right]$
$-\mathbb{E}\left[\langle\lambda_{1}(0),L_{0}\rangle+\int_{0}^{T}\langle\lambda_{1}^{1}(t),L_{t}^{1}-\nabla_{x}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t}\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)(L_{t}^{2}-\nabla_{x}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})L_{t})^{\top}\right)\mathrm{d}t\right$.
$+\left.\int_{0}^{T}\left\langle\lambda_{1}^{1}(t),\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right\rangle\mathrm{d}t+\int_{0}^{T}\mathrm{tr}\left(\lambda_{1}^{2}(t)\left(\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,\widehat{X}_{t},\widehat{\mu}_{t},\widehat{\alpha}_{t})(\widehat{X}_{t}^{\prime})L_{t}^{\prime}\right]\right)^{\top}\right)\mathrm{d}t\right]$
$\qquad+\left\langle\lambda_{2},\nu\right\rangle_{(C([0,T];\mathbb{R})^{**})^{n},(\mathcal{M}[0,T])^{n}},$ (22)
where $\lambda_{1}\in\mathbb{M}^{2}$ has the decomposition $\lambda_{1}(\cdot)=\lambda_{1}(0)+\int_{0}^{\cdot}\lambda_{1}^{1}(s)\mathrm{d}s+\int_{0}^{\cdot}\lambda_{1}^{2}(s)\mathrm{d}W_{s}$, $L\in\mathbb{T}^{2}$ has the representation $L_{\cdot}=L_{0}+\int_{0}^{\cdot}L_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}L_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}$, and $\nu$ is the associated radon measure of the mapping $\int_{0}^{\cdot}G(s)\mathrm{d}L_{s}^{3}\in\mathrm{BV}^{n}[0,T]$. Define an element $p=(p_{t})_{t\in[0,T]}$ in $\mathbb{M}^{2}$ that, for $t\in[0,T]$,

### v3

@@ 4 SMP by the Lagrange Multipliers Method
Let us first introduce some basic spaces of functions and processes. Let $\mathcal{X}$ be the space $L^{2}([0,T];\mathbb{R}^{l})/\mathord{\sim}$ equipped with the inner product $\langle\gamma^{1},\gamma^{2}\rangle_{\mathcal{X}}=\int_{0}^{T}\langle\gamma_{t}^{1},\gamma_{t}^{2}\rangle\mathrm{d}t$. Then, $(\mathcal{X},\langle\cdot,\cdot\rangle_{\mathcal{X}})$ is a Hilbert space. Let $\mathbb{L}_{\mathbb{F}}^{2,{l}}$ be the set of $\mathbb{F}$-adapted processes $\beta=(\beta_{t})_{t\in[0,T]}$ taking values in $\mathbb{R}^{l}$ such that $\mathbb{E}[\int_{0}^{T}|\beta_{t}|^{2}\mathrm{d}t]<\infty$. As a result, $\mathbb{L}_{\mathbb{F}}^{2,{l}}$ is a Hilbert space equipped with the inner product $\langle\beta^{1},\beta^{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}=\mathbb{E}[\int_{0}^{T}\langle\beta^{1}_{t},\beta^{2}_{t}\rangle\mathrm{d}t]$ for $\beta^{1},\beta^{2}\in\mathbb{L}_{\mathbb{F}}^{2,{l}}$. It can be easily verified that $\mathscr{U}[0,T]$ is closed and convex in $(\mathbb{L}_{\mathbb{F}}^{2,{l}},\langle\cdot,\cdot\rangle_{\mathbb{L}_{\mathbb{F}}^{2,{l}}})$. Define $\mathbb{M}^{2}$ ($\subset\mathbb{L}_{\mathbb{F}}^{2,{n}}$) as the set of $\mathbb{F}$-adapted (continuous) $\mathbb{R}^{n}$-valued process $\eta=(\eta_{t})_{t\in[0,T]}$ satisfying the Itô’s representation that
@@
where $\eta_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ is $\mathcal{F}_{0}$-measurable, $(\eta^{1},\eta^{2})=(\eta_{t}^{1},\eta_{t}^{2})_{t\in[0,T]}\in\mathbb{L}_{\mathbb{F}}^{2,n}\times\mathbb{L}_{\mathbb{F}}^{n\times r}$ take values respectively in $\mathbb{R}^{n}$ and $\mathbb{R}^{n\times r}$. Thus, $\mathbb{M}^{2}$ is a Hilbert space under the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ defined by
@@
To adopt the methodology introduced in [7], we propose to embed the state process into a suitable Banach space. Without loss of generality, we can assume that for any singular control $\zeta\in\mathcal{A}^{k}[0,T]$, the associated Radon measures $\mu_{i}$ for $i\in I$ of the trajectory $t\mapsto\zeta_{t}(\omega)$ is singular w.r.t. the Lebesgue measure. Otherwise, we can decompose $\mathrm{d}\zeta_{t}$ using the Lebesgue decomposition w.r.t. the Lebesgue measure and absorb the absolutely continuous (regular) term to the drift $b$. Indeed, suppose that $d\zeta_{t}$ admits the decomposition $d\zeta_{t}=\zeta_{t}^{\rm a}dt+d\zeta_{t}^{\rm s}$, where $d\zeta_{t}^{\rm s}$ is a non-decreasing process that is singular with respect to the Lebesgue measure on $[0,T]$. Then, by setting $\tilde{\alpha}=(\alpha,\zeta^{\rm a})$ as the new regular control and $\tilde{\zeta}=\zeta^{\rm s}$ as the new singular control, and by defining the new drift coefficient as $\tilde{b}(t,x,\mu,u,u^{\prime})=b(t,x,\mu,u)+G(t)u^{\prime}$, we can consider the problem where the singular control is singular with respect to the Lebesgue measure.
Consider the following space:
$\mathscr{X}$ $:=\Big\{X=(X_{t})_{t\in[0,T]};~X\text{ is }\mathbb{F}\text{-adapted process with $X_{0}\in L^{2}(\Omega;\mathbb{R}^{n})$ and }$
$\qquad~X_{\cdot}=X_{0}+\int_{0}^{\cdot}X_{s}^{1}\mathrm{d}s+\int_{0}^{\cdot}X_{s}^{2}\mathrm{d}W_{s}+\int_{0}^{\cdot}G(s)\mathrm{d}X_{s}^{3}$
$\qquad~\text{with}~(X^{1},X^{2},X^{3})\in\mathbb{L}_{\mathbb{F}}^{2,{n}}\times\mathbb{L}_{\mathbb{F}}^{2,{n\times r}}\times\mathcal{A}^{k}[0,T]\Big\}.$ (16)
@@
$\|X\|_{\mathscr{X}}:=\sqrt{\|X^{1}\|_{\mathbb{L}_{\mathbb{F}}^{2,{n}}}^{2}+\|X^{2}\|_{\mathbb{L}_{\mathbb{F}}^{2,{n\times r}}}^{2}}+\sqrt{\mathbb{E}\left[\left|\int_{0}^{T}G(s)\mathrm{d}X_{s}^{3}\right|^{2}\right]},~~\forall X\in\mathscr{X}.$
By the assumption that, for any $X\in\mathscr{X}$, its 3rd component $X^{3}$ of the decomposition in $\mathscr{X}$ is singular w.r.t. to the Lebesgue measure. We further define the set $\mathbb{G}$ and the norm $\|\cdot\|_{L^{2}(\Omega;\mathcal{M}([0,T])^{n})}$ by
$\mathbb{G}:=\Big\{S=(S_{t})_{t\in[0,T]};~S_{\cdot}=\int_{0}^{\cdot}G(s)\mathrm{d}\eta_{s}~\text{for}~\eta=(\eta_{t})_{t\in[0,T]}\in\mathcal{A}^{k}[0,T]$
[…]
$\|S\|_{L^{2}(\Omega;\mathcal{M}([0,T])^{n})}$ $:=\left\{\mathbb{E}\left[\left(\sum_{i=1}^{n}\|S^{i}\|_{[0,T];{\rm FM}}\right)^{2}\right]\right\}^{\frac{1}{2}}$
$=\left\{\mathbb{E}\left[\left|\sup_{\begin{subarray}{c}\ell:[0,T]\to\mathbb{R}^{n}\\ \|\ell^{i}\|_{\infty}+\mathrm{Lip}(\ell^{i})\leq 1\end{subarray}}\int_{[0,T]}\ell(t)^{\top}\mathrm{d}S_{t}\right|^{2}\right]\right\}^{\frac{1}{2}}.$
Here, the last equality follows by recalling the definition of the Foutet-Mourier norm. In view of (B6) of Assumption 2, we derive that $G(t)^{\top}G(t)$ is invertible for every $t\in[0,T]$. Therefore, for any mapping $h:[0,T]\to\mathbb{R}^{k}$ such that every component $h^{i}:[0,T]\times\mathbb{R}$ satisfies $\|h^{i}\|_{\infty}+\mathrm{Lip}(h^{i})\leq 1$, there exists a mapping $\ell:[0,T]\to\mathbb{R}^{n}$ such that $G(t)^{\top}\ell(t)=h(t)$ holds for all $t\in[0,T]$. Indeed, we can choose $\ell(t)=G(t)(G(t)^{\top}G(t))^{-1}h(t)$. By using the fact that $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$ and $|y^{\top}(G(t)^{\top}G(t))^{-1}y|\leq c^{-1}|y|^{2}$ for $y\in\mathbb{R}^{k}$, we deduce that $|\ell(t)-\ell(t)|\leq C|t-s|$ for some constant $C>0$ only depending on $c$. Consequently, it holds that
$\|\eta\|_{L^{2}(\Omega;\mathcal{M}([0,T])^{k})}^{2}$ $=\mathbb{E}\left[\left|\sup_{\begin{subarray}{c}h:[0,T]\to\mathbb{R}^{k}\\ \|h^{i}\|_{\infty}+\mathrm{Lip}(h^{i})\leq 1\end{subarray}}\int_{[0,T]}h(t)^{\top}\mathrm{d}\eta_{t}\right|^{2}\right]$
$=\mathbb{E}\left[\left|\sup_{\begin{subarray}{c}\ell:[0,T]\to\mathbb{R}^{n}\\ \|\ell^{i}\|_{\infty}+\mathrm{Lip}(\ell^{i})\leq 1\end{subarray}}\int_{[0,T]}\ell(t)^{\top}G(t)\mathrm{d}\eta_{t}\right|^{2}\right]\leq C\|S\|_{L^{2}(\Omega;\mathcal{M}([0,T])^{n})},$
where, $S(t)=\int_{0}^{t}G(t)\mathrm{d}\eta_{t}$ for $t\in[0,T]$ and the constant $C>0$ only depends on $(c,\|G\|_{\infty})$. This implies that $\mathbb{G}$ is a closed convex cone of the Banach space $L^{2}(\Omega;\mathcal{M}([0,T])^{n})$. Thus, $\mathscr{X}$ can be viewed as a closed convex cone of the Banach space $\mathbb{T}^{2}:=\mathbb{M}^{2}\times L^{2}(\Omega;({\cal M}[0,T])^{n})$ under the product norm. For any $\lambda=(\lambda^{1},\lambda^{2})\in\mathbb{T}^{2,*}\simeq\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{n})$, the dual pair of $X=L+S\in\mathscr{X}$ for $(L,S)\in\mathbb{M}^{2}\times\mathbb{G}^{2}$ is defined by
$\langle X,\lambda\rangle_{\mathbb{T}^{2},\mathbb{T}^{2,*}}=\langle L,\lambda^{1}\rangle_{\mathbb{M}^{2}}+\langle S,\lambda^{2}\rangle_{L^{2}(\Omega;({\cal M}[0,T])^{k}),L^{2}(\Omega;(C([0,T];\mathbb{R})^{**})^{k})}.$
@@
$\mathtt{X}:=\mathbb{T}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}\times L^{2}(\Omega;({\cal M}[0,T])^{n}),\quad\mathtt{Y}_{i}:=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},~\forall i\in I,\quad\mathtt{H}:=\mathbb{T}^{2}.$ (17)
@@
$J(X,\alpha,\zeta)$ $:=\mathbb{E}\left[\int_{0}^{T}f(t,X_{t},\mathcal{L}(X_{t}),\alpha_{t})\mathrm{d}t+{\int_{[0,T]}c(t)d\zeta_{t}+g(X_{T},\mu_{T})}\right],$
@@
$\begin{cases}\displaystyle\inf~J(X,\alpha,\zeta)\\[3.99994pt] \displaystyle\mathrm{s.t.}~~(X,\alpha,\zeta)\in C:=\mathscr{X}\times\mathscr{U}[0,T]\times\mathcal{A}^{k}[0,T],\\[3.99994pt] \displaystyle~~~~~~~\Phi_{i}(X,\alpha,\zeta)\leq_{K}0,~~\forall i\in I,\\[3.99994pt] \displaystyle~~~~~~~F(X,\alpha,\zeta)=0,\end{cases}$ (19)
@@
By Theorem 3.10 and Corollary 3.11, the existence of an optimal control is guaranteed. Moreover, as discussed in [7], the Fritz–John (FJ) condition provides a necessary condition for the optimal regular controls. In the sequel, we shall extend this methodology by introducing the FJ condition for C-MFC problem with singular controls. By using Lemma 3.11 in [7], the FJ optimality condition holds for the constrained optimization problem (19) in the following sense.
@@
Sppose that Assumption refass3 holds. Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal pair to problem (19). Then, there exist multipliers $r_{0}\geq 0$, $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ and $\lambda=(\lambda_{1},\lambda_{2})\in\mathbb{M}^{2}\times L^{2}(\Omega;(C([0,T],\mathbb{R})^{**})^{n})$, which are not all zeros, such that
@@
$\qquad\quad-\left\langle\lambda,D_{\alpha}F(\widehat{X},\widehat{\alpha},\widehat{\zeta})(\cdot)\right\rangle_{\mathtt{H}^{*},\mathtt{H}}~~\text{\rm in}~\mathbb{L}_{\mathbb{F}}^{2,{l}},$
@@
[…]
$N_{C}(\widehat{\boldsymbol{x}}):=\{\xi\in\mathtt{X}^{*}:\langle\xi,\widehat{\boldsymbol{x}}-\boldsymbol{x}\rangle_{\mathtt{X}^{*},\mathtt{X}}\geq 0,~\forall\boldsymbol{x}\in C\},\quad\widehat{\boldsymbol{x}}\in\mathtt{X},$ (21)
and $D_{\zeta}$ denotes the Fréchet derivative with respect to $\zeta$ in the Banach space $\mathcal{M}([0,T])^{k}$.
@@
Sppose that Assumption refass3 holds. Let $(\widehat{X},\widehat{\alpha},\widehat{\zeta})\in C$ be an optimal solution to problem (3). There exist some $r_{0}\geq 0$ and $\eta^{i}\in\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}+}$ for $i\in I$ such that
@@
Furthermore, the stochastic (first-order) minimum condition holds that, $\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.
@@
$Y_{t}$ $={r_{0}\nabla_{x}g(\widehat{X}_{T},\widehat{\mu}_{T})+r_{0}\mathbb{E}^{\prime}\left[\partial_{\mu}g(\widehat{X}_{T}^{\prime},\widehat{\mu}_{T})(\widehat{X}_{T})\right]}$
$\quad+\int_{t}^{T}\nabla_{x}H^{r_{0}}(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s},Y_{s},Z_{s})\mathrm{d}s+\int_{t}^{T}\mathbb{E}^{\prime}\left[\partial_{\mu}H^{r_{0}}(s,\widehat{X}_{s}^{\prime},\widehat{\mu}_{s},\widehat{\alpha}_{s}^{\prime},Y_{s}^{\prime},Z_{s}^{\prime})(\widehat{X}_{s})\right]\mathrm{d}s$
@@
$\langle\xi_{1},\widehat{X}-X\rangle_{\mathbb{T}^{2}}+\langle\xi_{2},\widehat{\alpha}-\alpha\rangle_{\mathbb{L}_{\mathbb{F}}^{2,{l}}}+\langle\xi_{3},\zeta-\widehat{\zeta}\rangle_{(C([0,T];\mathbb{R})^{**})^{n},({\cal M}[0,T])^{n}}\geq 0.$
Firstly, we claim that $r_{0}$ and $(\eta^{i})_{i\in I}$ cannot be zero simultaneously. Otherwise, we assume that $r_{0}=0$ and $\eta^{i}=0$ for $i\in I$ as the 1st step in the proof of Theorem 3.12 in [7]. Therefore, it holds by setting $\widehat{\alpha}=\alpha$, $\widehat{\zeta}=\zeta$ and Lemma 4.1 that, for every $X\in\mathscr{X}$,
$0$ $\leq\langle\xi_{1},\widehat{X}-X\rangle_{\mathtt{T}^{2}}=\langle\lambda,D_{X}F(\widehat{X},\widehat{\alpha})(X-\widehat{X})\rangle_{\mathtt{H}^{*},\mathtt{H}}$
$=X_{t}-\widehat{X}_{t}-\int_{0}^{t}\nabla_{x}b(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})^{\top}(X_{s}-\widehat{X}_{s})\mathrm{d}s-\int_{0}^{t}\nabla_{x}\sigma(s,\widehat{X}_{s},\widehat{\mu}_{s},\widehat{\alpha}_{s})^{\top}(X_{s}-\widehat{X}_{s})\mathrm{d}W_{s}$